# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v2

@@ 6.2. Decode Evaluation
Second, relative gains of FA_vAttention and FA_Paged increase over vLLM with the batch size e.g., as batch size increases from 4 to 32 for Llama-3-8B, relative gains increase from 1.05×1.05$\times$1.05 × to 1.58×1.58$\times$1.58 ×. This is because the latency of a decode attention kernel is proportional to the total number of tokens in the batch (vidur, ). Therefore, the contribution of attention kernel in the overall latency – and consequently gains with a more efficient kernel – increase with the batch size. Further, for the same reasons as discussed in §6.1 (faster attention and lower CPU overhead), FA_vAttention delivers up to 1.23×1.23$\times$1.23 × higher throughput than FI_Paged (Yi-6B, batch size 12).
@@
Finally, note that vAttention is only as good as the state-of-the-art PagedAttention for decode throughput, as compared to prefills where it outperforms PagedAttention. We believe this is due to memory bound nature of decode attention i.e., memory stalls make it possible to hide the effect of additional compute that paging support requires. This is evident from Table 6 wherein FA_Paged and FA_vAttention have similar latency. However, hiding compute overhead in a prefill attention kernel is hard because it is already compute bound (see Table 5 for latency gains in the prefill attention kernel).
6.3. End-to-end Throughput Evaluation
For end-to-end performance, we measure the makespan of serving a static trace of 50 requests for long context scenarios. To understand system performance for different workloads, we vary the initial context length of requests from 32K to 128K, and the ratio of prefill to decode tokens (referred as P:D) from 500 to 50. Note that a higher P:D as well as a longer context indicates that the workload is more prefill bound; we expect vAttention to improve performance in such cases.
Figure 8 shows the makespan for all the three models. For Yi-6B, FA_vAttention is up to 1.22×1.$2$, 1.16×1.16$\times$1.16 × and 1.13×1.13$\times$1.13 × faster than FA_Paged for P:D ratio of 500, 100 and 50. FA_vAttention is also up to 1.19×1.19$\times$1.19 × faster for Llama-3-8B and up to 1.16×1.16$\times$1.16 × faster for Yi-34B, compared to FA_Paged. Our gains over FA_Paged are higher at higher P:D ratios, and for a fixed P:D, our gains increase as the context length grows. These results are consistent with our expectation since a higher P:D and context lengths both mean that the system spends more time processing prefills where use of non-paged attention kernels outperform paged kernels.
Interestingly, we observe that despite being faster than FA_Paged in prefills, FI_Paged under-performs FA_Paged in end-to-end throughput in most cases (due to inefficient decodes as shown in Figure 7). Hence, our gains over FI_Paged are relatively higher (up to 1.29×1.29$\times$1.29 ×), except for some cases where FI_Paged outperforms FA_Paged (e.g., Yi-6B and Yi-34B, context length=128K, P:D=500).
@@ 6.4.1. Hiding allocation latency
Figure 9 shows the latency of more than 2500 decode iterations of Llama-3-8B. For this experiment, we used a batch of 32 requests and initialized the prefill context length of each request to be in between 4K–8K (chosen randomly). It is evident that overlapping memory allocation with compute effectively hides the latency of allocating memory via CUDA APIs. We used 2MB pages for this experiment to show that even the worst case latency can be hidden (Table 2 shows that 2MB pages incur highest latency). In contrast, allocating memory synchronously via CUDA APIs leads to frequent latency spikes of 5ms–20ms, depending on how many requests need new physical memory pages.
@@ 6.4.2. Deferred reclamation
Figure 10 shows that synchronous memory allocation using CUDA APIs for prefills incurs overhead of up to 1.15×1.15$\times$1.15 × with 64KB pages and up to 1.03×1.03$\times$1.03 × with 2MB pages. In most cases, deferred reclamation eliminates the need to invoke CUDA APIs for prefills because a newly arrived request can simply re-use physical memory pages that were allocated to a prior request. This way, deferred reclamation ensures that prefill latency is not affected by memory allocation.
6.4.3. Memory allocation bandwidth
Table 7 shows that even with 64KB (our smallest) page size, vAttention can allocate as much as 7.6GB per second per GPU. This is more than an order of magnitude higher than the maximum memory allocation rate of 600MB per second of decodes (Figure 4). Larger page sizes and higher TP dimensions increase the memory allocation rate proportionally. Therefore, memory allocation bandwidth of CUDA APIs is more than enough for LLM inference.
@@ 6.4.3. Memory allocation bandwidth
6.4.4. Effect of page size.
In many applications, use of smaller pages can potentially degrade performance due to TLB thrashing (neighbourhoodawareat, ; ingens, ; hawkeye, ; architecturalsupportforgpuat, ). We find that this is not the case for LLM inference. For example, Figure 11 shows that the execution latency of an attention kernel remains largely unaffected when the page size for KV-cache is reduced to 64KB, from the original 2MB. In a separate experiment, we find that these results are also consistent with very large models e.g., Llama-3-70B and GPT-3-175B. We attribute this to the regular computation pattern of the attention operator as well as the hand-tuned implementations that explicitly try to avoid irregular memory accesses.
Further, we find that 2MB pages are good enough for many online serving scenarios where latency constraints limit the maximum batch size. However, for throughput-oriented scenarios, smaller pages may be preferred. For example, using 2MB pages could serve batch sizes of up to 187, 203 and 56 for Yi-6B, Llama-3-8B and Yi-34B on a dynamic trace (dataset: OpenChat (wang2023openchat, ), load: 7 queries per seconds). In contrast, 64KB pages helped serve batch sizes of up to 240, 258 and 68 for our models (see Figure 12).
Figure 11. Effect of page size on the performance of FlashAttention’s prefill (left) and decode (right) attention kernels (model: Llama-3-8B).
Figure 12. Maximum batch size obtained with different page sizes for a dynamic workload. Large pages limit batch size due to internal fragmentation.
Figure 13. Illustration of code changes needed to replace the prefill attention kernel of FlashAttention by FlashInfer when using vAttention for dynamic physical memory allocation.
Model # Tokens in a physical memory block Max memory waste per request
64KB 128KB 256KB 2MB 64KB 128KB 256KB 2MB
Yi-6B (TP-1) 64 128 256 2048 4MB 8MB 16MB 128MB
Yi-6B (TP-2) 128 256 512 4096 8MB 16MB 32MB 256MB
[…]
Llama-3-8B (TP-2) 64 128 256 2048 8MB 16MB 32MB 256MB
[…]
Yi-34B (TP-2) 64 128 256 2048 15MB 30MB 60MB 480MB
Table 8. Block size (number of tokens in a physical memory page) as a function of the page size and degree of tensor parallelism. Columns on the right show how much physical memory can be wasted per-request in the worst-case.
@@ 6.5. Programming Effort
vAttention makes it feasible to replace one attention kernel with another with only a few lines of code changes in the serving framework (Figure 13). In contrast, in PagedAttention, a developer first needs to write a paged kernel and then make significant changes in the serving framework. For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes in over 15 files (vllmfi1, ; vllmfi2, ; vllmfi3, ). Implementing the initial paging support in FlashAttention kernel also required $\approx 280$ lines of code changes (fapaging, ) and additional efforts to enable support for smaller block sizes (fasmall, ).
7. Discussion
Batch Size FA_Paged (Block Size = 256) FA_Paged (Block Size = 16)
Model Ratio
Yi-6B 1 0.085 0.085 1.00×1.$0$
2 0.093 0.094 1.01×1.01$\times$1.01 ×
4 0.109 0.115 1.05×1.05$\times$1.05 ×
[…]
2 0.097 0.105 1.07×1.07$\times$1.07 ×
4 0.136 0.143 1.05×1.05$\times$1.05 ×
[…]
2 0.088 0.095 1.08×1.08$\times$1.08 ×
4 0.116 0.117 1.01×1.01$\times$1.01 ×
Table 9. Impact of varying block size on the execution latency (in milliseconds) of FlashAttention decode kernel.
7.1. Further Analysis of Internal Fragmentation
vAttention uses page size as the unit of physical memory allocation as opposed to the PagedAttention’s unit of KV-cache block size. Therefore, it is natural to ask what is the relationship between page size and block size. To compute this relationship, let $S$ be the size of a single token’s per-layer K-cache (or V-cache) on a worker (see §5.1). Then, page size t translates to KV-cache block size $=t/S$. Note that $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16). Therefore, block size = $t/(H\times D\times P)$.
Table 8 shows the block size for different model configurations and physical memory allocation sizes; we show each model under two TP configurations – TP-1 and TP-2 – to highlight the effect of TP dimension on block size. In addition, the table also shows how much physical memory can be (theoretically) wasted by vAttention due to internal fragmentation in the worst-case. The worst-case occurs when a new page is allocated but remains completely unused.
Context Length Batch Size FI_NonPaged FI_Paged Ratio
8K 4 0.902 0.071 12.7×12.7$\times$12.7 ×
8K 8 0.931 0.128 7.3×7.3$\times$7.3 ×
8K 16 0.943 0.229 4.2×4.2$\times$4.2 ×
[…]
16K 8 1.821 0.233 7.8×7.8$\times$7.8 ×
16K 16 1.875 0.450 4.2×4.2$\times$4.2 ×
[…]
32K 8 3.291 0.458 7.2×7.2$\times$7.2 ×
32K 16 3.365 0.891 3.8×3.8$\times$3.8 ×
Table 10. Latency (in milliseconds) of paged and non-paged decode attention kernels in FlashInfer. Ratio denotes the latency of non-paged kernel divided by the latency of paged kernel (model: Llama-3-8B, 2 A100 GPUs).
vAttention allocates physical memory equivalent to the page size on each TP worker whereas the per-token physical memory requirement of a worker goes down as TP dimension increases (because KV heads get split across TP workers). Therefore, block size increases proportionally with the TP dimension. Table 8 shows that using 2MB pages leads to large KV-cache block sizes of 1024 (Llama-3-8B and Yi-34B, TP-1) to 4096 (Yi-6B, TP-2) which could waste significant amount of physical memory (100s of MBs, per-request). Use of smaller (e.g., 64KB) pages reduces internal fragmentation proportionally, resulting in KV-cache block size of 32 (Yi-34B and Llama-3-8B, TP-1) to 128 (Yi-6B, TP-2), which in turn limit the maximum theoretical waste of only 4-15MB physical memory, per request. This shows that controlling the granularity of physical memory allocation is essential for reducing fragmentation. If required, page size can be reduced further to as low as 4KB which is the minimum page size supported in almost all architectures today, including NVIDIA GPUs.
7.2. Effect of Block Size on FlashAttention Kernel
We find that varying the block size of KV-cache also affects the execution latency of FlashAttention’s decode kernel (in addition to vLLM wherein using a larger block size increase the latency of decode kernel by up to 3×3$\times$3 × as shown in Figure 3). For example, Table 9 shows that using a smaller block size of 16 increases the latency of FlashAttention decode kernel by up to $5\%$ for Yi-6B, up to $8\%$ for Llama-3-8B and up to $9\%$ for Yi-34B. This is likely because work distribution and shared memory usage in a paged kernel needs to explicitly account for block size. Hence, writing an implementation that is equally efficient for all block sizes could be non-trivial. In fact, it is interesting to see that vLLM decode kernel performs better with small block sizes whereas FlashAttention decode kernel performs better with large block sizes. While the impact of block size on kernel latency is much less pronounced in FlashAttention compared to vLLM, a degradation in performance is still undesirable for an operator as important as attention. vAttention is free of such implementation challenges because it does not require a block table.
7.3. FlashInfer Non-paged Decode Attention Kernel
Note that we did not include FlashInfer non-paged decode kernel in our evaluation (§6). There are two reasons for this. First, the non-paged kernel has a much higher latency than the paged counterpart of FlashInfer e.g., Table 10 shows that the latency of non-paged kernel is up to 14.6×14.6$\times$14.6 × higher than the paged kernel. Further, the latency of the paged kernel follows an expected trend i.e., increasing the context length or batch size leads to a proportional increase in the execution latency. In contrast, increasing the batch size does not affect the latency of the non-paged kernel (e.g., context lengths of 8K and 32K). This behavior of the non-paged kernel is counter-intuitive and likely due to un-optimized implementation. Note that the job of a memory allocator is to simply allocate memory – it cannot accelerate a kernel. Second, the non-paged decode kernel of FlashInfer does not support variable sequence lengths – a fundamental requirement for LLM inference.
@@ 8. Related Work
Optimizing LLM inference is an active area of research. Various techniques have been proposed to improve different aspects of LLM serving like batching (orca, ; vllmsosp, ; sarathiserve2024, ), disaggregation (patel2023splitwise, ; distserve2024, ; tetriinfer, ), scheduling (flexgen, ; fastserve, ). However, central to all these techniques is the need for efficient KV-cache memory management. Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trtllmgithub, ), LightLLM (lightllm:github, ), and libraries e.g., FlashAttention (fagithub, ) and FlashInfer (figithub, ). In contrast, vAttention offers an alternate – which we believe is also a more principled – approach to dynamic KV-cache memory management.
In a recent work, GMLake (gmlakeasplos24, ) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size. In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors for large object allocations. In contrast, vAttention is focused on avoiding fragmentation for LLM inference. Different from training, LLM inference is latency sensitive and requires smaller granularity allocations. We proposed various LLM inference specific optimizations to meet these requirements.
Similar to our approach, PyTorch also recently added support for expandable segments (expandable-segments, ) that dynamically attach physical memory pages to a pre-reserved virtual memory buffer, if enabled. However, PyTorch uses expandable segments opportunistically, using synchronous memory allocation, and allocates physical memory pages only in multiples of 2MB granularity. In contrast, vAttention caters to the specific requirements of the KV-cache.
Some other concurrent works are also motivated to optimize attention kernels (flashattnhopper, ; mirage-tensor-compiler, ; thunderkittens, ; heavyhitters, ; flashdecoding++, ; leanattention, ; flash-attention-3, ). As one would expect, new research ideas do not start with paging support in mind. vAttention would make it easier to deploy them.

### v3

@@ 7.2. Decode Evaluation
Second, relative gains of FA2_vAttention and FA2_Paged increase over vLLM with the batch size, e.g., as batch size increases from 4 to 32 for Llama-3-8B, relative gains increase from 1.05×1.05$\times$1.05 × to 1.58×1.58$\times$1.58 ×. This is because the latency of a decode attention kernel is proportional to the total number of tokens in the batch (Agrawal et al., 2024a). Therefore, the contribution of attention kernel in the overall latency – hence gains with a more efficient kernel – increase with the batch size. Further, for the same reasons as discussed in §7.1 (faster attention and lower CPU overhead), FA2_vAttention delivers up to 1.23×1.23$\times$1.23 × higher throughput than FI_Paged (Yi-6B, batch size 12).
@@
Finally, note that vAttention is only as good as the state-of-the-art PagedAttention for decode throughput, as compared to prefills where it outperforms PagedAttention. This is because in case of decode attention, paged and non-paged kernels have similar latency as shown in Table 7. We believe this is due to memory bound nature of decode attention, i.e., memory stalls make it possible to hide the effect of additional compute that paging support requires. However, hiding compute overhead in a prefill attention kernel is hard because it is already compute bound (see Table 6 for vAttention’s gains in the prefill kernel).
Figure 9. Offline inference throughput while serving long-context requests from the arXiv-Summarization dataset (arx, 2024).
7.3. End-to-end Performance: Offline Scenarios
We measure the end-to-end system throughput in an offline scenario as the number of requests completed per minute. We evaluate a total of 427 long-context requests from the arXiv-Summarization workload trace wherein the total context length per-request varies from 64K tokens to 192K tokens and the number of output tokens per-request varies from 17 to 5153. The mean prefill to decode token ratio is 356. Figure 9 shows throughput for different attention backends for all three models. FA2_vAttention outperforms FA2_Paged by 1.18×1.18$\times$1.18 ×, 1.15×1.15$\times$1.15 × and 1.13×1.13$\times$1.13 ×, and FI_Paged by 1.19×1.19$\times$1.19 ×, 1.23×1.23$\times$1.23 ×, and 1.14×1.14$\times$1.14 × for Yi-6B, Llama-3-8B and Yi-34B, respectively. In general, vAttention’s performance gains are proportional to the context length and the prefill to decode token ratio (P:D ratio); these factors determine how much prefill attention contributes to the total runtime. A higher P:D ratio as well as longer context lengths indicate that the workload is more prefill bound. vAttention provides higher gains in such cases.
7.4. End-to-end Performance: Online Scenarios
We evaluate an online inference scenario serving long-context requests from arXiv-Summarization (arx, 2024). In total, we run 512 requests wherein the per-request input context length varies from 22K to 45K tokens (mean 29K), the number of decode tokens varies from 6 to 3250 (mean 348) and the mean P:D ratio is 129. We evaluate performance near the serving capacity of the system which denotes the maximum load a system can handle without incurring high queuing delays (Kwon et al., 2023; Agrawal et al., 2024b). We vary input load (queries-per-second or QPS) based on Poisson distribution and schedule requests in first-come-first-serve order. Figure 10 shows the CDF of the end-to-end request execution latency for this experiment. We see that vAttention consistently outperforms both baselines. For example, FA2_vAttention reduces the median request execution latency over FA2_Paged by up to 42% (QPS 0.25) for Yi-6B, by up to 28% (QPS 0.3) for Llama-3-8B and by up to 29% for Yi-34B (QPS 0.1). The primary reason for vAttention’s performance gain is that it can compute the prefill phase of new requests faster than the PagedAttention-based systems which substantially reduces queuing delays. Consistent with our prior results, FI_Paged is slower than FA2_Paged and hence our gains our FI_Paged are relatively higher compared to our gains over FA2_Paged.
Figure 10. CDF of end-to-end request execution latency in online inference under varying load.
Figure 11. Offline inference throughput on H100 GPUs.
7.5. Exemplifying Portability: FlashAttention-3
We demonstrate the portability benefit of vAttention with the recently released FA3 kernel (Shah et al., 2024). FA3 is optimized for the NVIDIA Hopper architecture and did not support PagedAttention when released. Therefore, dynamic memory allocation via PagedAttention is not feasible for FA3 at the time of writing this paper (integration of FA3 into vLLM is also a work in progress (fa3, 2024a, b)). vAttention not only enables dynamic memory allocation with FA3, it also requires no code changes to deploy FA3. Figure 11 shows the offline inference throughput of our models with 1–2 H100 GPUs (Yi-6B deployed on a single GPU and the other models on two GPUs each) on the same arXiv-Summarization-based workload as evaluated in §7.3. FA3 with vAttention provides an additional speedup of up to 1.35×1.35$\times$1.35 × (Yi-6B) over FA2_vAttention which is already up to 1.15×1.15$\times$1.15 × faster than FA2_Paged.
7.6. Ablation Studies
@@ 7.6.1. Hiding allocation latency
Figure 12 shows the latency of more than 500 decode iterations of Llama-3-8B. For this experiment, we used a batch of 32 requests with per request prefill context length varying between 2K to 3K tokens. It is evident that overlapping memory allocation with compute effectively hides the latency of CUDA VMM APIs. We used 2MB pages for this experiment to show that even the worst case memory allocation latency can be hidden by overlapping it with compute (Table 3 shows that 2MB pages incur highest latency). In contrast, allocating memory synchronously via CUDA APIs leads to frequent latency spikes of 5 to 15 milliseconds, depending on how many requests require physical memory allocation in a given iteration.
@@ 7.6.2. Deferred reclamation
Figure 13 shows that synchronous memory allocation incurs prefill overhead of up to 1.15×1.15$\times$1.15 × with 64KB pages and up to 1.03×1.03$\times$1.03 × with 2MB pages. In most cases, deferred reclamation eliminates the need to invoke CUDA VMM APIs for prefills because a newly arrived request can simply re-use physical memory allocated to a prior request. This way, deferred reclamation ensures that prefill latency is not affected by memory allocation.
@@
Figure 14. Effect of page size on the runtime of FlashAttention-2’s prefill (left) and decode (right) attention kernels (model: Llama-3-8B).
7.6.3. Effect of page size.
In many applications, use of smaller pages can potentially degrade performance due to TLB thrashing (Shin et al., 2018; Kwon et al., 2016; Panwar et al., 2019; Pichai et al., 2014). We find that this is not the case for LLM inference. For example, Figure 14 shows that the execution latency of an attention kernel remains largely unaffected when the KV cache is allocated using 64KB pages, compared to using 2MB pages. In a separate experiment, we find that these results are also consistent with very large models, e.g., Llama-3-70B and GPT-3-175B. We attribute this to the regular computation pattern of the attention operator as well as the hand-tuned implementations that explicitly try to avoid irregular memory accesses.
Model Block Size (# Tokens in a page-group)
64KB 128KB 256KB 2MB
Yi-6B (TP-1) 64 128 256 2048
Yi-6B (TP-2) 128 256 512 4096
Llama-3-8B (TP-1) 32 64 128 1024
Llama-3-8B (TP-2) 64 128 256 2048
Yi-34B (TP-1) 32 64 128 1024
Yi-34B (TP-2) 64 128 256 2048
Table 8. KV cache block size as a function of page-group size and degree of tensor parallelism.
Table 8 shows the block size i.e., number of tokens per physical page-group. Smaller page-groups enable fine-grained allocation, approaching vLLM’s recommended block size of 16–32 (the minimum block size in FlashAttention-2 is 256 – higher than vAttention). This shows that vAttention is as effective in reducing fragmentation as PagedAttention. In terms of end-to-end system performance, we find that 2MB pages are good enough for many online serving scenarios where latency constraints prevent use of very high batch sizes. In contrast, smaller page-groups are better for throughput-oriented scenarios where maximizing batch size is important to obtain peak performance. For example, using 2MB pages could serve batch sizes of up to 187, 203 and 56 for Yi-6B, Llama-3-8B and Yi-34B on a dynamic trace (dataset: OpenChat (Wang et al., 2023), load: 7 queries per seconds). In contrast, 64KB pages helped serve batch sizes of up to 240, 258 and 68 for our models as shown in Figure 15.
Figure 15. Maximum batch size obtained with different page-group sizes for a dynamic workload.
@@
7.6.4. Memory allocation bandwidth
Table 9 shows that even with 64KB pages (our smallest), vAttention can allocate as much as 7.6GB physical memory per second per GPU. This is more than an order of magnitude higher than the maximum memory allocation rate of 750MB per second of decodes (Figure 4). Larger page-group sizes and higher TP dimensions increase the memory allocation rate proportionally. Therefore, memory allocation bandwidth of CUDA VMM APIs is more than sufficient for LLM inference.
8. Discussion
8.1. Managing KV cache via Unified Memory
To enable dynamic memory allocation, we also considered leveraging unified memory via cudaMallocManaged (NVIDIA, 2024; Nayak et al., 2021). However, we find that unified memory support is currently not suitable for serving LLMs. First, it does not support partial freeing, preventing reclamation of physical memory of individual requests. Second, it lacks support for memory aliasing which prevents de-duplication of KV cache content in physical memory (de-duplication is useful when requests share a common prefix (Kwon et al., 2023)), consequently limiting batch size. cudaMallocManaged also allocates 2MB pages by default which can cause severe fragmentation. However, our changes in NVIDIA drivers are based on unified memory: we allocate virtual tensors using cudaMallocManaged, enabling support for partial freeing and page sharing with additional APIs. In addition, we introduce latency hiding optimizations and support for smaller pages. Therefore, one could consider our extensions to NVIDIA drivers as “unified memory optimized for LLM serving”.
Model Block size (# Tokens in a 2MB page)
w/o Tensor Slicing w/ Tensor Slicing
Yi-6B (TP-1) 2048 64
Yi-6B (TP-2) 4096 128
Llama-3-8B (TP-1) 1024 32
Llama-3-8B (TP-2) 2048 64
Yi-34B (TP-1) 1024 18
Yi-34B (TP-2) 2048 36
Table 10. KV cache block size with and without tensor slicing with 2MB pages.
8.2. Reducing Fragmentation via Tensor Slicing
To reduce fragmentation caused by 2MB pages, we also provide an alternative method that does not require modifying NVIDIA drivers. In this method, we use a single 2MB page to store the KV cache tokens of all layers for a given request. This can be done by allocating one virtual tensor of shape $[B,L,N,H,D]$ for K cache (and one for V cache) and slicing it across all layers, instead of allocating $2\times N$ virtual tensors of shape $[B,L,H,D]$. This reduces fragmentation to $1/N$ of the prior design as shown in Table 10. In this design, the K cache (or V cache) of a particular layer $n$ is represented by slice $[B,L,n:n+1,H,D]$ of the original tensor, and can be passed to attention kernels for computation. However, note that with tensor slicing, the K cache (and V cache) of each layer is no longer contiguous. Computing attention over such a tensor slice is possible only if the the attention kernel supports memory addressing with strides. While many kernels (e.g., FlashAttention-2) support strides out-of-the-box, the earlier versions of FlashInfer lacked such support (fi-, 2024). Therefore, relying on tensor slicing as the primary method of reducing fragmentation would have prevented vAttention from supporting FlashInfer kernels. Hence, to support unmodified FlashInfer kernels, we chose to add support for smaller pages in NVIDIA drivers. Importantly, our two solutions are compatible with each other; one could deploy smaller pages with tensor slicing to further reduce fragmentation, if required.
@@ 8.3. Programming Effort
PagedAttention requires significant programming effort. For example, integrating FlashInfer decode kernels in vLLM required more than 600 lines of code changes in over 15 files (vll, 2024c, e, d). Implementing the initial paging support in FlashAttention-2 kernel also required $\approx 280$ lines of code changes (fap, 2024) and additional efforts to support small block sizes (fas, 2024a). In contrast, vAttention makes it feasible to replace one attention kernel with another with only a few lines of code changes in the serving framework (Figure 16). Note that the example shown in Figure 16 has no interaction with memory management; this is precisely the goal of vAttention, i.e., when a slow kernel is replaced by a fast kernel, memory management should continue to work transparently.
Figure 16. Illustration of code changes needed to replace the prefill attention kernel of FlashAttention-2 by FlashInfer when using vAttention for memory allocation.
@@ 9. Related Work
Optimizing LLM inference is an active area of research  (Mei et al., 2024; Griggs et al., 2024; Miao et al., 2023; Zhu et al., 2024; Kossmann et al., 2024; Strati et al., 2024; Lee et al., 2024; Singhania et al., 2024; Zheng et al., 2024; Ma et al., 2024; Mahmud et al., 2024). Various techniques have been proposed to improve different aspects of LLM serving like batching (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024b), disaggregation (Patel et al., 2024; Zhong et al., 2024; Hu et al., 2024), scheduling (Sheng et al., 2023; Wu et al., 2023). However, central to all these techniques is the need for efficient KV cache memory management. Since vLLM, PagedAttention has been adopted in various serving frameworks e.g., TensorRT-LLM (trt, 2023), LightLLM (lig, 2023), and libraries e.g., FlashAttention-2 (fag, 2022) and FlashInfer (fig, 2023). Some other concurrent works have also proposed optimizations for attention kernels (Bikshandi and Shah, 2023; Wu et al., 2024; thu, 2024; Zhang et al., 2023b; Hong et al., 2024; Sanovar et al., 2024; Shah et al., 2024; Kamath et al., 2024). Deploying new kernels for inference is hard with PagedAttention. vAttention offers an alternate – which we believe is also a more principled – approach to dynamic KV cache memory management that makes it easier to deploy new kernels.
In a recent work, GMLake (Guo et al., 2024) showed that using CUDA virtual memory support can mitigate fragmentation in DNN training jobs, increasing training batch size. In particular, GMLake uses CUDA support to coalesce multiple smaller physical memory pages into a single virtually contiguous object that can prevent out-of-memory errors. In contrast, vAttention targets optimizing inference workloads.